# Classical and quantum chaotic synchronization in coupled dissipative time crystals

### v2

@@ Abstract
We investigate the dynamics of two coherently coupled dissipative time crystals. In the classical mean-field limit of infinite spin length, we identify a regime of chaotic synchronization, marked by a positive largest Lyapunov exponent and a Pearson correlation coefficient close to one. At the boundary of this regime, the Pearson coefficient varies abruptly, marking a crossover between staggered and uniform $z$-magnetization. To address finite-size quantum dynamics, we employ a quantum-trajectory approach and study the trajectory-resolved expectations of subsystem $z$-magnetizations. Their histograms over time and trajectory realizations exhibit maxima that undergo a staggered-to-uniform crossover analogous to the classical one. In analogy with the classical case, we interpret this behavior as quantum chaotic synchronization, with dissipative quantum chaos evidenced by the steady-state density matrix exhibiting Gaussian Unitary Ensemble statistics. The locations of the classical and quantum crossover are consistent with each other, reflecting quantum-classical correspondence despite the noncommutativity of the infinite-spin and infinite-time limits and the impact of entanglement, quantified via the entanglement entropy between subsystems.
@@ 1 Introduction
A related concept is that of time crystals, where a many-body system exhibits persistent oscillations in an order parameter, thus breaking continuous time-translation symmetry [32, 33]. In driven, unitary systems, this behavior appears as a collective subharmonic response [34, 35], while in dissipative systems it emerges as spontaneous oscillations of collective magnetization in the thermodynamic limit [36]. In this limit, such behavior often reduces to mean-field nonlinear dynamics with self-sustained oscillations. This naturally raises the question of whether coupled time crystals can synchronize. Following the first investigation of synchronization in many dissipative time crystals [37], subsequent works refined this picture [38], culminating in a recent study of synchronization between two dissipative (or boundary) time crystals in the thermodynamic limit [39]. These results revealed distinct synchronized phases, with oscillations that can be either in phase or out of phase.
@@
To probe quantum effects at finite spin magnitude, we employ the quantum-trajectory method and analyze trajectory-resolved $z$-magnetizations. Evaluating for each subsystem the histogram of the $z$-magnetization expectations, as they change over time and trajectory realizations, one can see a maximum that displays a sharp crossover from a staggered to a uniform regime across the two subsystems. This behavior closely echoes the classical case and suggests the emergence of quantum chaotic synchronization. This interpretation is strengthened by the steady-state density matrix, whose spectrum exhibits Gaussian unitary ensemble statistics that marks quantum chaos, and by the behavior of the bipartite entanglement entropy, highlighting the central role of entanglement in quantum synchronization [13, 4, 5, 40]. Notably, the crossover points in the quantum and classical cases coincide, despite the noncommutativity of the infinite-time and infinite-spin limits: While the classical system exhibits persistent oscillations, the quantum dynamics relaxes to a nonequilibrium steady state [41]. Nevertheless, both regimes display a qualitatively and quantitatively similar synchronization crossover, although the physics is different and in the quantum case a role in the crossover is played by entanglement.
@@ 2.1 Lindblad equation and classical mean field limit
We consider a system consisting of two coherently coupled spins of magnitude $S$, evolving under a Lindblad dynamics as introduced in Ref. [39]. The unitary part of the evolution is provided by the following Hamiltonian (we will work in natural units, where $\hbar=1$),
@@
where $\hat{S}_{\alpha}^{j}$ can be considered as the sum of NN spin-$1/2$ degrees of freedom, $\hat{S}_{\alpha}^{j}=\frac{1}{2}\sum_{l=1}^{N}\hat{\sigma}_{j\,l}^{\alpha}$ (with $\alpha\in\{x,y,z\}$ and $j\in\{A,B\}$). Moreover, $\hat{S}_{j}^{\pm}=\hat{S}_{j}^{x}\pm i\hat{S}_{j}^{y}$. By choosing $S=N/2$, the system is restricted to the Hilbert subspace with maximum total spin of the NN spin-$1/2$ degrees of freedom. Therefore, the limit $S\to\infty$ corresponds to the thermodynamic limit, as shown in Ref. [36]. The parameter $\Gamma$ determines the strength of the coherent coupling between the subsystems and is renormalized by the factor $1/S$ to ensure the correct, extensive scaling of the interaction term in the thermodynamic limit. The parameter $\Omega$ is the driving frequency of the two subsystems when uncoupled ($\Gamma=0$).
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{B}^{z}}_{t}$ $=\Omega\braket{\hat{S}_{B}^{y}}_{t}+\frac{\Gamma}{2S}\left(\hat{S}_{B}^{y}\hat{S}_{A}^{x}-\hat{S}_{B}^{x}\hat{S}_{A}^{y}\right)+\frac{\kappa}{S}\left[(\hat{S}_{B}^{x})^{2}+(\hat{S}_{B}^{y})^{2}\right]\,$,
@@ 2.2 Quantum-trajectory approach for the finite-size case
We implement this dynamics using the quantum-jump unraveling of the Lindblad equation Eq. (2). One can show that the Lindblad evolution is equivalent to the average over infinite realizations of a pure-state dynamics with a stochastic non-Hermitian Schrödinger equation. For each Lindblad equations there are many possible stochastic-dynamics approaches leading to it and are called unravelings (see Refs. [42, 43, 44] for a review). Here we choose the so-called quantum-jump unraveling that takes the following form. We discretize the time dividing it in intervals $\delta t$ and in each time interval:
@@ 3.1 Methods
$\Lambda_{L}=\frac{\delta t}{T_{\rm max}}\sum_{n=1}^{T_{\rm max}/\delta t}\ln\left(\frac{d_{n}}{d_{0}}\right)\,.$
@@ 3.2 Results
Figure 1: Largest Lyapunov exponent $\Lambda_{L}$ versus $\Gamma$ and $\Omega$ for initial conditions $m_{j}^{z}=1$ and $m_{j}^{x,y}=0$ for all $j$ (panel a), averaged over random initial conditions as given in Sec. 3.1.2 with $a=0.1$ (panel b), average over random initial conditions as given in Sec. 3.1.2 with $a=1$ (panel c). The averages are over $N_{\rm r}=50$ initial conditions. See the main text and Fig. 2 for a discussion about the symbols shown in panel (b).
@@
Figure 2: Largest Lyapunov exponent $\braket{\Lambda_{L}}$ (top row) and Pearson correlation coefficient $\overline{C_{P}}$ (bottom row) versus $\Gamma$ along the line $\Gamma+\Omega=1.5$ (left column), $\Gamma+\Omega/2=2$ (center column), and $\Gamma+\Omega/2=1.5$ (right column). These values are calculated as averages over $N_{\rm r}=100$ initial conditions generated on a spherical cap with parameter $a=0.1$.
@@
We begin by plotting the LLE defined in Sec. 3.1.3, shown in Fig. 1. In panel (a) we choose a specific initial condition (all spins pointing along the north pole), while in panel (b) and (c) we average over initial conditions as given in Sec. 3.1.2, choosing respectively $a=0.1$ and $a=1$. All the regimes described in Ref. [39] for Fig. 1(a) are clearly visible also when we average over many initial conditions, though they appear smeared out. In particular, at the bottom left there is a regime of negative LLE (absence of chaos) corresponding to relaxation of the two magnetizations to a fixed point (trivial phase). Then there is a central lobe where the LLE is vanishing that corresponds to a continuous time crystal phase (marked as CTC3 in Ref. [39]). Another similar lobe appears in the upper left region, where there is another continuous time crystal phase (CTC1) which encompasses the usual dissipative time crystal (occurring when the coupling $\Gamma$ vanishes).
@@
Figure 3: Heat maps of various quantities versus $\Gamma$ and $\Omega$. (a,b) Pearson correlation coefficient. (c,d) Difference between the magnetizations averaged over time and initial conditions, $\langle\overline{\delta m}\rangle=\braket{\overline{m_{B}^{z}}}-\braket{\overline{m_{A}^{z}}}$. (e,f) Relative sign $s=\operatorname{sign}(\braket{\overline{m_{z}^{A}}}\braket{\overline{m_{z}^{B}}})$ of the two magnetizations. The calculations are performed by averaging over $N_{\rm r}=100$ random initial conditions generated on a spherical cap with parameter $a=0.1$ (first row), and $a=1$ (second row).
@@
In Fig. 1(b), neglecting the trivial phase, we consider three lines cutting the parameter space in Fig. 1(b) from the upper left to the bottom right part of the panel; These lines are $\Omega+\Gamma=1.5$, $\Gamma+\Omega/2=2$, and $\Gamma+\Omega/2=1.5$ and are marked in Fig. 1(b) with different symbols (blue circles, green triangles, and orange squares, respectively). Along these lines, we plot, using the same symbols, the LLE (Fig. 2, top row) and the corresponding average Pearson coefficient described in Sec. 3.1.4 (Fig. 2, bottom row) versus $\Gamma$. Both the LLE and the Pearson coefficient are evaluated over a set of random initial conditions with $a=0.1$. We see that there is a threshold value $\Gamma\simeq 1.2$ in panel 2(a), $\Gamma\simeq 1.6$ in panel 2(b), $\Gamma\simeq 1.3$ in panel 2(c)] where the LLE moves from zero to positive, coinciding with a sudden increase of the Pearson coefficient. At this threshold, the Pearson coefficient shifts from strongly negative to strongly positive values. Below this threshold the system is in the CTC3 phase: Here, the Lyapunov exponent is vanishing, and the negative Pearson coefficient indicates that the two subsystem oscillate out of phase with respect to each other. Above the threshold the system becomes chaotic – as marked by the Lyapunov exponent becoming positive – but the two oscillators behave in a strictly correlated way, as witnessed by the positive value of the Pearson coefficient. Therefore, in this regime, the two subsystems display full chaotic synchronization.
@@
The behaviors of the Lyapunov exponent and Pearson coefficient closely relate to the behavior of the magnetization averaged over time and initial-condition $m_{j}^{z}(t)$ ($j=A,B$). We define this average as $\braket{\overline{m_{j}^{z}}}$ and in Fig. 4 we plot it versus $\Gamma$ along the three lines considered above. We identify three regimes: (i) a small-$\Gamma$ regime, where the average magnetization is vanishing for both subsystems; (ii) an intermediate-$\Gamma$ regime, where the average magnetizations of the two subsystems have opposite sign (staggered magnetization regime); (iii) a large-$\Gamma$ regime, where the average magnetizations of the two subsystems are nonvanishing and nearly equal (uniform magnetization regime). We find that the crossover between the staggered and uniform magnetization regimes coincides with the onset of full chaotic synchronization, as we can see by comparing the top row of Fig. 4 (where initial conditions are with $a=0.1$) with Fig. 2.
Figure 4: Time and initial-condition average values of $m_{j}^{z}(t)$, $\braket{\overline{m_{j}^{z}}}$ versus $\Gamma$, along the line $\Omega+\Gamma=1.5$ (a, left column), $\Gamma+\Omega/2=2$ (b, center column), and $\Gamma+\Omega/2=1.5$ (c, right column). We average over a set of $N_{\rm r}=100$ initial conditions generated either for $a=0.1$ (top row) or $a=1$ (bottom row). The legend is shared.
@@ 4 Quantum finite-size dynamics
Figure 5: (a,c) $\overline{S_{A}^{z}}$ and $\overline{S_{B}^{z}}$ versus $\Omega$ for different NN. (b,d) Time-averaged connected correlator $\overline{C_{AB}^{zz}}$. Average over $N_{\rm r}\geq 48$ trajectories. On the left column we consider the curves for the line $2\Omega+\Gamma=2$ and on the right one $2\Omega+\Gamma=1.5$.
@@
Further insight is gained by analyzing histograms of these quantities over time $t$ and across different trajectory realizations $r$ (as done in Refs. [48, 49]). We plot the histograms for $(S_{A}^{z})^{r}(t)=\braket{\psi_{t}^{r}|\hat{S}_{A}^{z}|\psi_{t}^{r}}$ and $(S_{B}^{z})^{r}(t)=\braket{\psi_{t}^{r}|\hat{S}_{B}^{z}|\psi_{t}^{r}}$ in Fig. 6(a,b). They are obtained as an histogram over a set of values obtained by changing $r$ and $t$. Along each trajectory, we also consider the entanglement entropy between the two subsystems, defined as
@@
Figure 7: (a,c) Maximum points of the histograms of $S_{A}^{z}$ and $S_{B}^{z}$ along and across trajectories versus $\Gamma$. (b,d) Maximum point of the histogram of the entanglement entropy versus $\Gamma$. We fix the constraint $2\Omega+\Gamma=2$ in the left column and $2\Omega+\Gamma=1.5$ in the right column. We take also $\delta t=10^{-4}$.
Figure 8: Heat map of $s=\operatorname{sign}(\braket{\overline{m_{A}^{z}}}\braket{\overline{m_{B}^{z}}})$ as shown in Fig. 3(f) where the lines $2\Omega+\Gamma=2$ (upper dotted line) and $2\Omega+\Gamma=1.5$ (lower dashed line) are drawn.
We notice that the threshold for this crossover in the quantum case is consistent with the mean field case, as we can see comparing Fig. 7(a,c) with Fig. 8. This figure is the same as Fig. 3(f) and depicts $s=\operatorname{sign}(\braket{\overline{m_{A}^{z}}}\braket{\overline{m_{B}^{z}}})$ that is -1 in case of staggered magnetization and +1 in case of uniform magnetization. On this plot we have drawn the lines $2\Omega+\Gamma=2$ (upper dotted line) and $2\Omega+\Gamma=1.5$ (lower dotted-dashed line). We see that the upper line crosses the classical boundary between staggered and uniform magnetization for $\Gamma\simeq 1.25$ that is consistent with the value where uniform magnetization sets in the corresponding quantum case [see Fig. 7(a)]. The same occurs for the lower line that crosses the boundary at $\Gamma\simeq 1$, consistently with the corresponding quantum case shown in Fig. 7(c). So we can see a classical-quantum correspondence in the threshold between staggered and uniform magnetization.
This happens although the infinite-$S$ limit and the infinite time-limit do not commute. Indeed, the infinite-$S$ limit shows persistent (periodic or aperiodic) oscillations, while the finite-$S$ size case relaxes for $t\to\infty$ to a NESS density matrix for any arbitrary large but finite $S$ [39]. Here we are studying the properties of this NESS density matrix, which are in general different from the ones of the long-time behavior in the mean-field regime (see examples in Refs. [52, 41]). So it is remarkable is that the crossover is present in both scenarios, and occurs at the same point, despite the non-commutativity of the limits and their different long-time behaviors. Additionally, entanglement plays a significant role in the quantum case but is absent classically. We also note that the noncommutative nature of these limits helps explain why the largest Lyapunov exponent (LLE), shown in Fig. 2, exhibits intervals of zero (regular dynamics) and positive (chaotic dynamics), whereas the NESS density matrix displays signatures of quantum chaos throughout the same parameter range (see Fig. 9). This difference arises from the LLE assessing the long-time dynamics of the mean-field, infinite-$S$ limit, whereas the NESS characterizes the long-time limit of finite-size quantum dynamics – the infinite-$S$ limit being taken first in the former case.

### v4

@@ Abstract
We investigate the dynamics of two coherently coupled dissipative time crystals. In the classical mean-field limit of infinite spin length, we identify a regime of chaotic synchronization, marked by a positive largest Lyapunov exponent and a Pearson correlation coefficient close to one. At the boundary of this regime, the Pearson coefficient varies abruptly, marking a crossover between staggered and uniform $z$-magnetization. To address finite-size quantum dynamics, we employ a quantum-trajectory approach and study the trajectory-resolved expectations of subsystem $z$-magnetizations. Their histograms over time and trajectory realizations exhibit maxima that undergo a staggered-to-uniform crossover analogous to the classical one. In analogy with the classical case, we interpret this behavior as quantum chaotic synchronization, with dissipative quantum chaos highlighted by the steady-state density matrix exhibiting Gaussian Unitary Ensemble statistics. The classical and quantum crossover points are different due to the noncommutativity of the infinite-time and infinite-spin-magnitude limits and the role played by entanglement in the quantum case, quantified via the two-subsystem entanglement entropy.
@@ 1 Introduction
A related concept is that of time crystals, where a many-body system exhibits persistent oscillations in an order parameter, thus breaking continuous time-translation symmetry [32, 33]. In driven, unitary systems, this behavior appears as a collective subharmonic response [34, 35], while in dissipative systems it emerges as spontaneous oscillations of collective magnetization in the thermodynamic limit [36]. In this limit, such behavior often reduces to mean-field nonlinear dynamics with self-sustained oscillations. This naturally raises the question of whether coupled time crystals can synchronize. Following the first investigation of synchronization in many dissipative time crystals [37], subsequent works refined this picture [38], culminating in a recent study of synchronization between two dissipative (or boundary) time crystals in the thermodynamic limit [39]. These results revealed distinct synchronized phases, with oscillations that can be either in phase or out of phase. (The dynamics of two coupled boundary time crystals has also been considered from a different perspective in [40, 41].)
@@
To probe quantum effects at finite spin magnitude, we employ the quantum-trajectory method and analyze trajectory-resolved $z$-magnetizations. Evaluating for each subsystem the histogram of the $z$-magnetization expectations, over time and the ensemble of trajectory realizations, one can see a maximum that displays a sharp crossover from a staggered to a uniform regime across the two subsystems. This behavior closely echoes the classical case and suggests the emergence of quantum chaotic synchronization. This interpretation is strengthened by the steady-state density matrix, whose spectrum exhibits Gaussian unitary ensemble statistics that marks quantum chaos, and by the behavior of the bipartite entanglement entropy, highlighting the central role of entanglement in quantum synchronization [13, 4, 5, 42]. Notably, the crossover points in the quantum and classical cases do not coincide, due to the noncommutativity of the infinite-time and infinite-spin limits: While the classical system exhibits persistent oscillations, the quantum dynamics relaxes to a nonequilibrium steady state [43]. Nevertheless, both regimes display a qualitatively similar synchronization crossover, although the physics is different and in the quantum case a role in the crossover is played by entanglement.
@@ 2.1 Lindblad equation and classical mean field limit
We consider a system consisting of two coherently coupled spins of magnitude $S$, evolving under the Lindblad dynamics introduced in Ref. [39]. The unitary part of the evolution is provided by the following Hamiltonian (we will work in natural units, where $\hbar=1$),
@@
where $\hat{S}_{\alpha}^{j}$ can be considered as the sum of NN spin-$1/2$ degrees of freedom, $\hat{S}_{\alpha}^{j}=\frac{1}{2}\sum_{l=1}^{N}\hat{\sigma}_{j\,l}^{\alpha}$ (with $\alpha\in\{x,y,z\}$ and $j\in\{A,B\}$). Moreover, $\hat{S}_{j}^{\pm}=\hat{S}_{j}^{x}\pm i\hat{S}_{j}^{y}$. By choosing $S=N/2$, the system is restricted to the Hilbert subspace with maximum total spin of the NN spin-$1/2$ degrees of freedom. Therefore, the limit $S\to\infty$ corresponds to a thermodynamic limit, as shown in Ref. [36]. The parameter $\Gamma$ determines the strength of the coherent coupling between the subsystems and is renormalized by the factor $1/S$ to ensure the correct, extensive scaling of the interaction term in the thermodynamic limit. The parameter $\Omega$ is the driving frequency of the two subsystems when uncoupled ($\Gamma=0$).
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\braket{\hat{S}_{B}^{z}}_{t}$ $=\Omega\braket{\hat{S}_{B}^{y}}_{t}+\frac{\Gamma}{2S}\left(\hat{S}_{B}^{y}\hat{S}_{A}^{x}-\hat{S}_{B}^{x}\hat{S}_{A}^{y}\right)+\frac{\kappa}{S}\left[(\hat{S}_{B}^{x})^{2}+(\hat{S}_{B}^{y})^{2}\right]\,.$
@@ 2.2 Quantum-trajectory approach for the finite-size case
We implement this dynamics using the quantum-jump unraveling of the Lindblad equation Eq. (2). One can show that the Lindblad evolution is equivalent to the average over infinite realizations of a pure-state dynamics with a stochastic non-Hermitian Schrödinger equation. For each Lindblad equations there are many possible stochastic-dynamics approaches leading to it and are called unravelings (see Refs. [44, 45, 46] for a review). Here we choose the so-called quantum-jump unraveling that takes the following form. We discretize the time dividing it in intervals $\delta t$ and in each time interval we proceed as follows.
@@ 3.1 Methods
$\Lambda_{L}=\frac{\delta t}{T_{\rm max}}\sum_{n=1}^{T_{\rm max}/\delta t}\ln\left(\frac{d_{n}}{d_{0}}\right)\,.$ (17)
@@ 3.2 Results
Figure 1: Largest Lyapunov exponent $\Lambda_{L}$ versus $\Gamma$ and $\Omega$ for initial conditions $m_{j}^{z}=1$ and $m_{j}^{x,y}=0$ for all $j$ (panel a), averaged over random initial conditions as given in Sec. 3.1.2 with $a=0.1$ (panel b), and with $a=1$ (panel c). The averages are over $N_{\rm r}=50$ initial conditions. See the main text and Fig. 2 for a discussion about the symbols shown in panel (b).
@@
Figure 2: Largest Lyapunov exponent $\braket{\Lambda_{L}}$ (top row) and Pearson correlation coefficient $\overline{C_{P}}$ (bottom row) versus $\Gamma$ along the line $\Gamma+\Omega=1.5$ (left column), $2\Omega+\Gamma=2$ (center column), and $2\Omega+\Gamma=1.5$ (right column). These values are calculated as averages over $N_{\rm r}=100$ initial conditions generated on a spherical cap with parameter $a=0.1$.
@@
We begin by plotting the LLE defined in Sec. 3.1.3, shown in Fig. 1. In panel (a) we choose a specific initial condition (all spins pointing along the north pole), while in panels (b) and (c) we average over initial conditions as given in Sec. 3.1.2, choosing respectively $a=0.1$ and $a=1$. All the regimes described in Ref. [39] for Fig. 1(a) are clearly visible also when we average over many initial conditions, though they appear smeared out. In particular, at the bottom left there is a regime of negative LLE (absence of chaos) corresponding to relaxation of the two magnetizations to a fixed point (trivial phase). Then there is a central lobe where the LLE is vanishing (again, absence of chaos) that corresponds to a continuous time crystal phase (marked as CTC3 in Ref. [39]). Another similar lobe appears in the upper left region, where there is another continuous time crystal phase (CTC1) which encompasses the usual dissipative time crystal (occurring when the coupling $\Gamma$ vanishes).
@@
Figure 3: Heat maps of various quantities versus $\Gamma$ and $\Omega$. (a,b) Pearson correlation coefficient. (c,d) Difference between the magnetizations averaged over time and initial conditions, $\langle\overline{\delta m}\rangle=\braket{\overline{m_{B}^{z}}}-\braket{\overline{m_{A}^{z}}}$. (e,f) Relative sign $s=\operatorname{sign}(\braket{\overline{m_{z}^{A}}}\braket{\overline{m_{z}^{B}}})$ of the two magnetizations. The calculations are performed by averaging over $N_{\rm r}=100$ random initial conditions generated on a spherical cap with parameter $a=0.1$ (first row), and $a=1$ (second row). The lines in panel (f) are discussed in Sec. 4.
@@
In Fig. 1(b), we consider three lines cutting the parameter space in Fig. 1(b) from the upper left to the bottom right part of the panel; These lines are $\Omega+\Gamma=1.5$, $2\Omega+\Gamma=2$, and $2\Omega+\Gamma=1.5$ and are marked in Fig. 1(b) with different symbols (blue circles, green triangles, and orange squares, respectively). Along these lines, we plot, using the same symbols, the LLE (Fig. 2, top row) and the corresponding average Pearson coefficient described in Sec. 3.1.4 (Fig. 2, bottom row) versus $\Gamma$. Both the LLE and the Pearson coefficient are evaluated over a set of random initial conditions with $a=0.1$. First of all, we notice in panels 2(c,d) that there is a regime of negative Lyapunov exponent corresponding to the trivial phase. Beyond that, we see that there is a threshold value $\Gamma\simeq 1.2$ in panel 2(a), $\Gamma\simeq 1.25$ in panel 2(b), $\Gamma\simeq 1.15$ in panel 2(c)] where the LLE moves from zero to positive, coinciding with a sudden increase of the Pearson coefficient. At this threshold, the Pearson coefficient shifts from strongly negative to strongly positive values. Below this threshold the system is in the CTC3 phase: Here, the Lyapunov exponent is vanishing, and the negative Pearson coefficient indicates that the two subsystem oscillate out of phase with respect to each other. Above the threshold the system becomes chaotic – as marked by the Lyapunov exponent becoming positive – but the two oscillators behave in a strictly correlated way, as witnessed by the positive value of the Pearson coefficient. Therefore, in this regime, the two subsystems display full chaotic synchronization.
@@
Figure 4: Time and initial-condition average values of $m_{j}^{z}(t)$, $\braket{\overline{m_{j}^{z}}}$ versus $\Gamma$, along the parameter-space line $\Omega+\Gamma=1.5$ (a, left column), $2\Omega+\Gamma=2$ (b, center column), and $2\Omega+\Gamma=1.5$ (c, right column). We average over a set of $N_{\rm r}=100$ initial conditions generated either for $a=0.1$ (top row) or $a=1$ (bottom row). The legend is shared. The background shadowing marks the Lyapunov exponent with the same color code as Fig. 1.
The behaviors of the Lyapunov exponent and Pearson coefficient closely relate to the behavior of the magnetization averaged over time and initial conditions. We define this average as $\braket{\overline{m_{j}^{z}}}$ ($j=A,B$) and in Fig. 4 we plot it versus $\Gamma$ along the three parameter-space lines considered above. We identify many regimes. In panel 4(a) we see (i) a small-$\Gamma$ regime, where the average magnetization is vanishing for both subsystems; (ii) an intermediate-$\Gamma$ regime, where the average magnetizations of the two subsystems have opposite sign (staggered magnetization regime); (iii) a large-$\Gamma$ regime, where the average magnetizations of the two subsystems are nonvanishing and nearly equal (uniform magnetization regime). We find that regime (i) corresponds to the CTC1 phase, regime (ii) to the CTC3 regime and regime (iii) to the chaotic synchronized phase, as highlighted by the shadow scale in the background marking the value of the Lyapunov exponent. (Note that CTC1 and CTC3 are separated by a small chaotic region.) In particular, the crossover between the regime (ii) of staggered magnetization and regime (iii) of uniform magnetization coincides with the onset of full chaotic synchronization, as we can see by comparing the top row of Fig. 4 (where initial conditions are with $a=0.1$) with Fig. 2. We can see regimes (ii) and (iii) also in panels 4(b,c), and also here the crossover between them corresponds to the onset of chaotic synchronization, as further highlighted by the shadowing that marks the value of the Lyapunov exponent. Here, regime (i) is absent, because there is no CTC1 phase along these lines in the parameter space, but there is the trivial phase instead. Here there is a staggered magnetization and the crossover to CTC3 is marked by a discontinuity in the derivative, closely mirroring the behavior of the Lyapunov exponent.
@@ 4 Quantum finite-size dynamics
Figure 5: (a,c) $\overline{S_{A}^{z}}$ and $\overline{S_{B}^{z}}$ versus $\Omega$ for different NN. (b,d) Time-averaged connected correlator $\overline{C_{AB}^{zz}}$. Average over $N_{\rm r}\geq 48$ trajectories. On the left column we consider the curves for the line $2\Omega+\Gamma=2$ and on the right one $2\Omega+\Gamma=1.5$. The vertical line marks the classical staggered-uniform crossover as found in Fig. 3(f).
@@
We see here that there is no correspondence between the quantum and the classical behavior, highlighted by the vertical line that marks the value of the staggered-uniform crossover along the considered parameter-space line, as read in Fig. 3(f).
Further insight is gained by analyzing histograms of these quantities over time $t$ and over the ensemble of different trajectory realizations $r$ (as done in Refs. [50, 51]). We plot the histograms for $(S_{A}^{z})^{r}(t)=\braket{\psi_{t}^{r}|\hat{S}_{A}^{z}|\psi_{t}^{r}}$ and $(S_{B}^{z})^{r}(t)=\braket{\psi_{t}^{r}|\hat{S}_{B}^{z}|\psi_{t}^{r}}$ in Fig. 6(a,b). They are obtained as an histogram over a set of values obtained by changing $r$ and $t$. Along each trajectory, we also consider the entanglement entropy between the two subsystems, defined as
@@
We notice that the threshold for this crossover in the quantum case is in general not consistent with the mean field case, as we can see comparing Fig. 7(a,c) with Fig. 3(f). We see this fact comparing the curves with the vertical line in the plots that marks the classical crossover from the staggered-magnetization to the uniform-magnetization regime, as read from 3(f). While along the parameter-space line $2\Omega+\Gamma=2$ there is a reasonable correspondence (left column of Fig. 7), this is not true along a different parameter-space line ($2\Omega+\Gamma=1.5$ – left column of Fig. 7. Nevertheless, the quantum behavior is qualitatively similar to the classical one, because in both Fig. 4 and in Fig. 6 one sees a similar staggered-uniform crossover.
@@
Figure 7: (a,c) Maximum points of the histograms of $S_{A}^{z}$ and $S_{B}^{z}$ along and across trajectories versus $\Gamma$. (b,d) Maximum point of the histogram of the entanglement entropy versus $\Gamma$. We fix the constraint $2\Omega+\Gamma=2$ in the left column and $2\Omega+\Gamma=1.5$ in the right column. We take also $\delta t=10^{-4}$. The vertical line marks the classical staggered-uniform crossover as found in Fig. 3(f).
The fact that the quantum and the classical crossover do not coincide is related with the noncommutativity of the infinite-$S$ limit and the infinite time-limit. Indeed, the infinite-$S$ limit shows persistent (periodic or aperiodic) oscillations, while the finite-$S$ size case relaxes for $t\to\infty$ to a NESS density matrix for any arbitrary large but finite $S$ [39]. Here we are studying the properties of this NESS density matrix, which are in general different from the ones of the long-time behavior in the mean-field regime (see examples in Refs. [54, 43]). So it is remarkable that the crossover is present in both scenarios, despite the non-commutativity of the limits and their different long-time behaviors. Additionally, entanglement plays a significant role in the quantum case but is absent classically. We also note that the noncommutative nature of these limits helps explain why the largest Lyapunov exponent, shown in Fig. 2, exhibits intervals of zero (regular dynamics) and positive (chaotic dynamics), whereas the NESS density matrix displays signatures of quantum chaos throughout the same parameter range (see Fig. 8). This difference arises from the LLE assessing the long-time dynamics of the mean-field, infinite-$S$ limit, whereas the NESS characterizes the long-time limit of finite-size quantum dynamics – the infinite-$S$ limit being taken first in the former case.